# Constraints on the Variation of the QCD Interaction Scale $\Lambda_{\text{QCD}}$

### v3

@@ Abstract
Introduction
The search for the variation of fundamental constants is a well-established area of research motivated by theories beyond the Standard model that unify gravity with the other fundamental forces, as well as cosmological models involving scalar fields which predict a space-time variation of these constants. Constraints on such variation have been derived from Big Bang Nucleosynthesis (BBN), quasar absorption spectra, the cosmic microwave background (CMB) and the Oklo natural nuclear reactor, see e.g. the review [1] and references therein. The stringent bounds on the linear drift of the fundamental constants have been obtained using measurements of the temporal drift of atomic clock transition frequencies. These experiments measure the variation of the dimensionless ratio of two transition frequencies $\nu_{1}$ and $\nu_{2}$
@@ Introduction
which have different dependencies on the fundamental constants. As a result, these measurements may be used to place constraints on the variation of these constants. Experimental results have been obtained by measuring the oscillating frequency ratios of electron transitions in a range of systems, including Dy/Cs [2], Rb/Cs [3], Yb/Cs [4], Sr/H/Si cavity [5], Cs/cavity [6], Cs/H [7], Al+/Hg+ [8] and Yb+/Yb+/Sr [9, 10]. The results of these experiments have been used to place constraints on the variations of fundamental constants using the sensitivity calculations performed in Refs. [11, 12, 13, 14, 15, 16, 17] and several other papers referenced there.
@@
It is important to note that if the absolute measurement of a transition frequency is performed, the dimensionless ratio of this frequency to the Cs hyperfine transition frequency, which defines the units of frequency and time, is being measured. The Cs transition frequency has a complicated dependence on the fundamental constants [15]. This principle applies to any measurement of variations of fundamental constants in any field of physics and astrophysics. Any measurement of a dimensionful parameter is, in fact, a measurement of the dimensionless ratio of that parameter to the unit used to measure it, which itself depends on the fundamental constants.
@@
Because any measurable quantity is ultimately a dimensionless ratio, it can always be written solely in terms of dimensionless fundamental constants. To do so, one first identifies the particular dimensionless constants that enter a given experiment. By contrast, many papers discuss variations of dimensionful parameters, such as the speed of light $c$, Planck’s constant $\hbar$, Newton’s gravitational constant $G_{N}$, the QCD scale $\Lambda_{\text{QCD}}$, or particle masses. The problem is that the numerical values of these quantities depend on the system of units: in natural units, for example, $c=1$ by definition, so it cannot ”vary.” Any apparent change can be absorbed by a re-definition of units. Consequently, interpreting experimental results in terms of changing dimensionful constants is not model-independent; one must instead identify a set of dimensionless fundamental constants relevant to this measurement and calculate the dependence of an observed effect on these constants  [^1] It was noted in Refs. [18, 19], that the number of possible ”dimensionless” constants may increase when we allow for fast variations of the constants, where ”fast” is determined by the time scale of the response of the studied species or experimental apparatus used. In this case, the relevant dimensionless quantity is, for example, the ratio $m_{e}/\expectationvalue{m_{e}}$, and $\expectationvalue{m_{e}}$ is the time average. In this sense, one may say that the experimental signal depends on the variation of dimensionful constants (the electron mass $m_{e}$ in this example). Sensitivity to such variations requires two systems, one which has a faster response (such as an atom) and another which is more inertial (such as a cavity).
@@
Alternatively, one may speak about the apparent variation of the fundamental constants, which may be just a convenient way to calculate the effects of specific terms in the interaction of Standard model particles with a dark matter (or dark energy [20, 21, 22, 23, 1]) scalar or pseudoscalar field $\phi$. For example, the interaction of fermions with $\phi$, which may be presented in the form $g\phi^{n}m\bar{f}f$, is similar to the fermion mass term $m\bar{f}f$. This means that one may define a variable fermion effective mass $m_{eff}=m(1+g\phi^{n})$, which depends on the field $\phi$. Here, $n=1,2$ for scalars and $n=2$ for pseudoscalars. Similarly, the interaction of $\phi$ with the electromagnetic field may be incorporated as a variable fine structure constant $\alpha_{eff}=\alpha(1+g\phi^{n})$. The effects of the variation of fermion masses and the variation of $\alpha$ have been calculated; see, e.g. [11, 12, 13, 24, 25, 14, 15, 16, 17]. Note that in this approach, the variation of measurement units is still a relevant issue if one considers measurements of dimensionful parameters. We note that using experimental results to place constraints on the variation of dimensionful parameters is only possible within a specific model of variation, e.g. an interaction with an ultralight dark matter field or dark energy field, which varies on a cosmological scale.
@@
We seek to examine the sensitivity of various systems to variations of the QCD scale $\Lambda_{\text{QCD}}$. Such a sensitivity appears in the frequency comparison of two microwave atomic clocks or optical and microwave clocks (see below). It is usually assumed that the comparison of two optical clock transitions is sensitive to variations of the fine structure constant only; however, following Ref. [9], we argue that this system is also sensitive to variations of $\Lambda_{\text{QCD}}$ and repurpose existing results to place constraints. We also examine the sensitivity of the fractional variation of the transition frequency between the ground state and the low-energy isomer in 229Th to variations in the QCD scale $\Lambda_{\text{QCD}}$, which appears to be enhanced relative to the proposed nuclear clock’s sensitivity to the quark mass $m_{q}$. Additionally, we place constraints on $\Lambda_{\text{QCD}}$ from quasar absorption spectra and Big Bang Nucleosynthesis. Finally, we reinterpret existing constraints from the Oklo natural nuclear reactor to place constraints on the variation of $\Lambda_{\text{QCD}}$.
I Constraints on $\Lambda_{\text{QCD}}$ via the Reinterpretation of Clock constraints
@@
where $X_{q,e}$ corresponds to the ratio of the quark/electron mass to the QCD scale $\Lambda_{\text{QCD}}$. Here, we are working in natural units, setting $c=1$ and assuming that $\Lambda_{\text{QCD}}$ has dimension of mass. For example, the variation of the ratio of two atomic, molecular or nuclear transition frequencies may be presented as
@@
Table 1 lists the current constraints on the variation of the ratio of the two hyperfine transition frequencies in 87Rb and 133Cs and the constraints on the variation in the ratio of optical transition frequencies in 1H, 199Hg+, 171Yb+ and 87Sr to the hyperfine transition in 133Cs. In the previous calculations in Refs. [14, 15], the dependence on $X_{q}$ was calculated assuming that $\Lambda_{\text{QCD}}$ is constant and only the quark masses vary. Below, we will demonstrate that there is no need for such an assumption.
The dependence of the atomic hyperfine transition frequency on $X_{q}$ appears mainly from the dependence on the nuclear magnetic moment. The dependence of the nuclear magnetic moments on $m_{q}$ was calculated in Refs. [14, 15]. A less significant contribution to the $m_{q}$ dependence, which originates from the variation of the nuclear radius, was calculated in Refs. [26, 27].
The nuclear magnetic moment and nuclear radius are also sensitive to the QCD scale $\Lambda_{\text{QCD}}$. In effective field theory, the nucleon magnetic moments receive loop-level corrections from pion-nucleon interactions governed by the spontaneously broken chiral symmetry of QCD. These corrections provide a natural mechanism by which these magnetic moments can acquire a dependence on both the quark mass and $\Lambda_{\text{QCD}}$. The nucleon magnetic moment including such corrections can be schematically written as [28, 29]:
@@
where $\mu_{N}^{(0)}$ denotes the static low-energy constant contributions at tree-level and $\mu_{N,\pi}^{(1)}$ represents the dominant corrections to the nucleon magnetic moments, measured in units of the nuclear magneton $e\hbar/2m_{p}c$, which at leading one-loop order are proportional to the pion mass $m_{\pi}$ [30, 31]:
@@
where $g_{A}$ is the axial coupling, $m_{N}$ the nucleon mass, and $f_{\pi}$ the pion decay constant. The sign of this correction is positive for the proton magnetic moment and negative for the neutron magnetic moment. Since $m_{\pi}^{2}\sim m_{q}\Lambda_{\text{QCD}}$ via the Gell-Mann–Oakes–Renner relation  [32] and $f_{\pi}\propto\Lambda_{\text{QCD}}$, the pion mass inherits a dependence on the light quark mass $m_{q}$ and $\Lambda_{\text{QCD}}$. Consequently, the one-loop correction scales as:
@@
87Rb 133 Cs $(-0.5\pm 5.3)\times 10^{-16}$ [33] $\alpha^{-0.49}X_{q}^{-0.025}$
1H 133Cs $(-32\pm 63)\times 10^{-16}$ [34] $\alpha^{2.83}X_{q}^{-0.039}X_{e}$
199Hg+ 133Cs $(3.7\pm 3.9)\times 10^{-16}$ [35] $\alpha^{6.05}X_{q}^{-0.039}X_{e}$
171Yb+ 133Cs $(-3.1\pm 3.4)\times 10^{-17}$ [36] $\alpha^{1.93}X_{q}^{-0.039}X_{e}$
87Sr 133Cs $(-1.0\pm 1.8)\times 10^{-15}$ [37] $\alpha^{2.77}X_{q}^{-0.039}X_{e}$
@@
Under the minimal assumption that this gluonic channel is the sole $\phi$ interaction, the constraints presented in Table 1 can be translated into limits on the temporal variation of the QCD scale, $\Lambda_{\text{QCD}}$, via the dependence of each transition on $X_{q}$ and $X_{e}$. The dominant contribution to the sensitivity arises from the $X_{e}$ term, and the tightest constraints may be obtained from the 171Yb+/Cs clock measurement [36]:
@@
I.1 Sensitivity of the comparison of two optical clock transition frequencies to $X_{e}$

### v4

@@ Abstract
I Introduction
The search for the variation of fundamental constants is a well-established area of research motivated by theories beyond the Standard model that unify gravity with the other fundamental forces, as well as cosmological models involving scalar fields which predict a space-time variation of these constants. Constraints on such variation have been derived from Big Bang Nucleosynthesis (BBN), quasar absorption spectra, the cosmic microwave background (CMB) and the Oklo natural nuclear reactor, see e.g. the review [Uzan] and references therein. The stringent bounds on the linear drift of the fundamental constants have been obtained using measurements of the temporal drift of atomic clock transition frequencies. These experiments measure the variation of the dimensionless ratio of two transition frequencies $\nu_{1}$ and $\nu_{2}$
@@ I Introduction
which have different dependencies on the fundamental constants. As a result, these measurements may be used to place constraints on the variation of these constants. Experimental results have been obtained by measuring the oscillating frequency ratios of electron transitions in a range of systems, including Dy/Cs [DyCs], Rb/Cs [Hees2016], Yb/Cs [YbCs], Sr/H/Si cavity [HSi], Cs/cavity [Tretiak], Cs/H [Fischer2004], Al+/Hg+ [AlHgdrift] and Yb+/Yb+/Sr [Banerjee2023, Filzinger2023]. The results of these experiments have been used to place constraints on the variations of fundamental constants using the sensitivity calculations performed in Refs. [PRLWebb, PRAWebb, CanJPh, Thomas, Tedesco, csquarks, Borschevsky] and several other papers referenced there.
@@
It is important to note that if the absolute measurement of a transition frequency is performed, the dimensionless ratio of this frequency to the Cs hyperfine transition frequency, which defines the units of frequency and time, is being measured. The Cs transition frequency has a complicated dependence on the fundamental constants [Tedesco]. This principle applies to any measurement of variations of fundamental constants in any field of physics and astrophysics. Any measurement of a dimensionful parameter is, in fact, a measurement of the dimensionless ratio of that parameter to the unit used to measure it, which itself depends on the fundamental constants.
@@
Because any measurable quantity is ultimately a dimensionless ratio, it can always be written solely in terms of dimensionless fundamental constants. To do so, one first identifies the particular dimensionless constants that enter a given experiment. By contrast, many papers discuss variations of dimensionful parameters, such as the speed of light $c$, Planck’s constant $\hbar$, Newton’s gravitational constant $G_{N}$, the QCD scale $\Lambda_{\text{QCD}}$, or particle masses. The problem is that the numerical values of these quantities depend on the system of units: in natural units, for example, $c=1$ by definition, so it cannot ”vary.” Any apparent change can be absorbed by a re-definition of units. Consequently, interpreting experimental results in terms of changing dimensionful constants is not model-independent; one must instead identify a set of dimensionless fundamental constants relevant to this measurement and calculate the dependence of an observed effect on these constants  [^1] It was noted in Refs. [Antypas1, Antypas2], that the number of possible ”dimensionless” constants may increase when we allow for fast variations of the constants, where ”fast” is determined by the time scale of the response of the studied species or experimental apparatus used. In this case, the relevant dimensionless quantity is, for example, the ratio $m_{e}/\expectationvalue{m_{e}}$, and $\expectationvalue{m_{e}}$ is the time average. In this sense, one may say that the experimental signal depends on the variation of dimensionful constants (the electron mass $m_{e}$ in this example). Sensitivity to such variations requires two systems, one which has a faster response (such as an atom) and another which is more inertial (such as a cavity).
@@
Alternatively, one may speak about the apparent variation of the fundamental constants, which may be just a convenient way to calculate the effects of specific terms in the interaction of Standard model particles with a dark matter (or dark energy [BarrowMagueijo1998, Barrow2005, BarrowSandvikMagueijo2002, Wetterich2003, Uzan]) scalar or pseudoscalar field $\phi$. For example, the interaction of fermions with $\phi$, which may be presented in the form $g\phi^{n}m\bar{f}f$, is similar to the fermion mass term $m\bar{f}f$. This means that one may define a variable fermion effective mass $m_{eff}=m(1+g\phi^{n})$, which depends on the field $\phi$. Here, $n=1,2$ for scalars and $n=2$ for pseudoscalars. Similarly, the interaction of $\phi$ with the electromagnetic field may be incorporated as a variable fine structure constant $\alpha_{eff}=\alpha(1+g\phi^{n})$. The effects of the variation of fermion masses and the variation of $\alpha$ have been calculated; see, e.g. [PRLWebb, PRAWebb, CanJPh, Shuryak2002, FlambaumShuryak, Thomas, Tedesco, csquarks, Borschevsky]. Note that in this approach, the variation of measurement units is still a relevant issue if one considers measurements of dimensionful parameters. We note that using experimental results to place constraints on the variation of dimensionful parameters is only possible within a specific model of variation, e.g. an interaction with an ultralight dark matter field or dark energy field, which varies on a cosmological scale.
@@
We seek to examine the sensitivity of various systems to variations of the QCD scale $\Lambda_{\text{QCD}}$. Such a sensitivity appears in the frequency comparison of two microwave atomic clocks or optical and microwave clocks (see below). It is usually assumed that the comparison of two optical clock transitions is sensitive to variations of the fine structure constant only; however, following Ref. [Banerjee2023], we argue that this system is also sensitive to variations of $\Lambda_{\text{QCD}}$ and repurpose existing results to place constraints. We also examine the sensitivity of the fractional variation of the transition frequency between the ground state and the low-energy isomer in 229Th to variations in the QCD scale $\Lambda_{\text{QCD}}$, which appears to be enhanced relative to the proposed nuclear clock’s sensitivity to the quark mass $m_{q}$. Additionally, we place constraints on $\Lambda_{\text{QCD}}$ from quasar absorption spectra and Big Bang Nucleosynthesis. Finally, we reinterpret existing constraints from the Oklo natural nuclear reactor to place constraints on the variation of $\Lambda_{\text{QCD}}$.
II Constraints on $\Lambda_{\text{QCD}}$ from atomic clocks measurements
II.1 Ratios of magnetic dipole hyperfine and optical transition frequencies
@@
where $X_{q,e}$ corresponds to the ratio of the quark/electron mass to the QCD scale $\Lambda_{\text{QCD}}$. Here, we are working in natural units, $c=\hbar=1$. For example, the variation of the ratio of two atomic, molecular or nuclear transition frequencies may be presented as
@@
Table 1 lists the current constraints on the variation of the ratio of the two hyperfine transition frequencies in 87Rb and 133Cs and the constraints on the variation in the ratio of optical transition frequencies in 1H, 199Hg+, 171Yb+ and 87Sr to the hyperfine transition in 133Cs. In the previous calculations in Refs. [Thomas, Tedesco], the dependence on $X_{q}$ was calculated assuming that $\Lambda_{\text{QCD}}$ is constant and only the quark masses vary. Below, we will demonstrate that there is no need for such an assumption.
The dependence of the atomic hyperfine transition frequency on $X_{q}$ appears mainly from the dependence on the nuclear magnetic moment. The dependence of the nuclear magnetic moments on $m_{q}$ was calculated in Refs. [Thomas]. A less significant contribution to the $m_{q}$ dependence, which originates from the variation of the nuclear radius, was calculated in Refs. [Wiringa2009, Dinh2009].
The nuclear magnetic moment and nuclear radius are also sensitive to the QCD scale $\Lambda_{\text{QCD}}$. In effective field theory, the nucleon magnetic moments receive loop-level corrections from pion-nucleon interactions governed by the spontaneously broken chiral symmetry of QCD. These corrections provide a natural mechanism by which these magnetic moments can acquire a dependence on both the quark mass and $\Lambda_{\text{QCD}}$. The nucleon magnetic moment including such corrections can be schematically written as [Bernard1995, Kubis2001]:
@@
where $\mu_{N}^{(0)}$ denotes the static low-energy constant contributions at tree-level and $\mu_{N,\pi}^{(1)}$ represents the dominant corrections to the nucleon magnetic moments, measured in units of the nuclear magneton $e\hbar/2m_{p}c$, which at leading one-loop order are proportional to the pion mass $m_{\pi}$ [Bernard1992, Jenkins1993]:
@@
where $g_{A}$ is the axial coupling, $m_{N}$ the nucleon mass, and $f_{\pi}$ the pion decay constant. The sign of this correction is positive for the proton magnetic moment and negative for the neutron magnetic moment. Since $m_{\pi}^{2}\sim m_{q}\Lambda_{\text{QCD}}$ via the Gell-Mann–Oakes–Renner relation [GellMan1968] and $f_{\pi}\propto\Lambda_{\text{QCD}}$, the pion mass inherits a dependence on the light quark mass $m_{q}$ and $\Lambda_{\text{QCD}}$. Consequently, the one-loop correction scales as:
@@
87Rb 133 Cs $(-1.4\pm 0.9)\times 10^{-16}$ [RbCsNew] $\alpha^{-0.49}X_{q}^{-0.025}$
1H 133Cs $(-32\pm 63)\times 10^{-16}$ [FischerHCs2004] $\alpha^{2.83}X_{q}^{-0.039}X_{e}$
199Hg+ 133Cs $(3.7\pm 3.9)\times 10^{-16}$ [Fortier2007] $\alpha^{6.05}X_{q}^{-0.039}X_{e}$
171Yb+ 133Cs $(-3.1\pm 3.4)\times 10^{-17}$ [Lange2021] $\alpha^{1.93}X_{q}^{-0.039}X_{e}$
87Sr 133Cs $(-1.0\pm 1.8)\times 10^{-15}$ [Blatt2008] $\alpha^{2.77}X_{q}^{-0.039}X_{e}$
@@
Under the minimal assumption that this gluonic channel is the sole $\phi$ interaction, the constraints presented in Table 1 can be translated into limits on the temporal variation of the QCD scale, $\Lambda_{\text{QCD}}$, via the dependence of each transition on $X_{q}$ and $X_{e}$. The dominant contribution to the sensitivity arises from the $X_{e}$ term, and the tightest constraints may be obtained from the 171Yb+/Cs clock measurement [Lange2021]:
@@
II.2 Ratio of electric quadrupole hyperfine to magnetic dipole hyperfine and optical transition frequencies
In our paper [FlambaumMansour2023] we found that the electric quadrupole hyperfine constant $B$ has higher sensitivity to variation of $m_{q}$ than the magnetic hyperfine constant $A$, and suggested to measure variation of the ratio $B/A$. In this subsection we will show that this ratio $B/A$, as well as the ratio of $B$ to an optical transition frequency, are highly sensitive to the variation of $\Lambda_{\text{QCD}}$.
The dependence of the hyperfine constants $B$ and $A$ on the fundamental constants is given by:
$B$ $\propto\frac{eQ}{a_{B}^{3}}R_{B}(Z\alpha)\sim\frac{e^{2}r_{0}^{2}}{a_{B}^{3}}R_{B}(Z\alpha)\,$, (8)
$A$ $\propto\frac{\mu_{B}\mu_{N}}{a_{B}^{3}}R_{A}(Z\alpha)\sim\frac{e^{2}\hbar^{2}g_{N}}{m_{e}m_{N}c^{2}a_{B}^{3}}R_{A}(Z\alpha)\,$, (9)
where $r_{0}\approx 1.2\ \text{fm}$ is the internucleon distance, $a_{B}$ is the Bohr radius, $\mu_{B}$ and $\mu_{N}$ are the Bohr and nuclear magnetons and $g_{N}$ is the nuclear magnetic $g$-factor, $Q$ is the electric quadrupole moment, $R_{B}(Z\alpha)$ and $R_{A}(Z\alpha)$ are the relativistic factors for the electric quadrupole and magnetic dipole constants respectively, which are presented e.g. in the paper [Tedesco] and the book [Sobelman]. Therefore, the ratio of the electric quadrupole and magnetic dipole constants may be written as
$\frac{B}{A}\propto\frac{r_{0}^{2}m_{e}m_{N}c^{2}R_{B}(Z\alpha)}{g_{N}\hbar^{2}R_{A}(Z\alpha)}\,.$ (10)
The dominating effect in the variation of this ratio comes from the variation of $r_{0}^{2}$ and $m_{N}$. The dependence of $g_{N}$ on hadronic parameters is generally weak and varies from nucleus to nucleus [Thomas, Tedesco]. To leading order, this dependence can be neglected, allowing for general estimates that are valid across all nuclei.
The variation of the internucleon distance $r_{0}$ in terms of the pion and nucleon mass has been calculated in Ref. [Wiringa2009]  [^2] The dependence on the nucleon and pion masses arises mainly from the kinetic energy and the long-distance behavior of the strong interaction. These contributions can be calculated reliably and are included in the present work. Ref. [Wiringa2009] also considered vector meson and $\Delta$ contributions, which describe short-distance nucleon interaction. The description of the short- distance interaction is model-dependent. Also, the $V$ and $\Delta$ contributions partially cancel. Since their inclusion does not clearly improve the accuracy, we omit the $V$ and $\Delta$ contributions.:
$\frac{\delta r_{0}}{r_{0}}=1.8\frac{\delta m_{\pi}}{m_{\pi}}-4.8\frac{\delta m_{N}}{m_{N}}\,.$ (11)
The variation of the nucleon mass $m_{N}$ can be related to the variation of the fundamental constants by [Thomas, Damour2010, SHIFMAN1978, Junnarkar2013, Aoki2020, Oswald2022]
[…]
where $m_{q,-}=(m_{u}-m_{d})$ is the mass difference. The dependence of the variation of the nucleon mass on this parameter is weak, and we may neglect it. We will also assume that $\frac{\delta m_{s}}{m_{s}}=\frac{\delta m_{q}}{m_{q}}$, where $m_{s}$ is the mass of the strange quark (this assumption may be motivated by the Higgs mechanism of mass generation). Noting the dependence of the pion mass on $\Lambda_{\text{QCD}}$ and $m_{q}$ from the Gell-Mann-Oakes-Renner relation [GellMan1968]
$m_{\pi}^{2}\sim\Lambda_{\text{QCD}}m_{q}\,$, (13)
we yield the following dependence on the hadronic parameters
$\frac{\delta r_{0}}{r_{0}}=-3.5\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.3\frac{\delta m_{q}}{m_{q}}\,.$ (14)
Thus, using Eqs. (10) and (14), we yield
$\frac{\delta(B/A)}{(B/A)}=-6.0\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.7\frac{\delta m_{q}}{m_{q}}\,.$ (15)
It may also be useful to estimate the dependence of the ratio of the electric quadrupole constant $B$ to the optical transition energy $E_{\text{opt}}\propto e^{2}/a_{B}$. Using Eq. (8) we obtain
$\frac{\delta(B/E_{\text{opt}})}{(B/E_{opt})}\approx\frac{\delta(r_{0}^{2}/a_{B}^{2})}{r_{0}^{2}/a_{B}^{2}}.$ (16)
We omit here the relativistic factors which contribute to the $\alpha$-dependence. The dependence on hadronic parameters may be presented as
$\frac{\delta(B/E_{\text{opt}})}{(B/E_{opt})}=-9.6\frac{\delta m_{N}}{m_{N}}+3.6\frac{\delta m_{\pi}}{m_{\pi}}$ (17)
$=-6.9\frac{\delta\Lambda_{\text{QCD}}}{\Lambda_{\text{QCD}}}+0.6\frac{\delta m_{q}}{m_{q}}.$ (18)
II.3 Sensitivity of the comparison of two optical clock transition frequencies to hadronic parameters